# The low multipoles in the Pantheon+SH0ES data

### v1

@@ Abstract
In previous work we have shown that the dipole in the low redshift supernovae of the Pantheon+SH0ES data does not agree with the one inferred from the velocity of the solar system as obtained from CMB data. We interpreted this as the presence of significant bulk velocities. In this paper we study the monopole, dipole and quadrupole in the Pantheon+SH0ES data. We find that in addition to the dipole also both, the monopole and the quadrupole are detected with high significance. They are of similar amplitudes as the bulk flow. While the monopole is only significant at very low redshift, the quadrupole even increases with redshift.
@@ 1 Introduction
Due to our motion with respect to the surface of last scattering, the CMB also exhibits a dipole with an amplitude of about $10^{-3}$. This dipole has been discovered in the 1970s [5, 6] and is now measured with exquisite precision [7, 8, 1]. This anisotropy in the CMB also leads to a correlation of adjacent multipoles which have consistently been measured with a significance of about 5 standard deviations [9]. Attributing the entire CMB dipole to our motion, one infers a velocity of the solar system given by
@@
where $({\rm ra,dec})$ are the ‘right ascension’ (ra) and ‘declination’ (dec) denoting the directions with respect to the baricenter of the solar system (at J2000, i.e. January 1, 2000). A possible intrinsic dipole in the CMB of the same order as the higher multipoles is expected to change this result by about 1%.
@@
Within the standard model of cosmology we expect to see this dipole due to our motion also in the large scale distribution of galaxies [10]. While first results of a radio survey agreed reasonably well with the CMB velocity [11], more recent analyses of catalogs of radio galaxies and quasars have found widely differing results from which significantly larger peculiar velocities have been inferred [12, 13, 14, 15, 16]. The latest results [17] show a $5\sigma$ discrepancy with the CMB dipole which is considered by the authors as a challenge of the cosmological principle. There are, however also critiques that the analysis of the data might be too simplified and that a more refined analysis would give results that are consistent with standard cosmology [18, 19, 20], see also [21] for an alternative method which gives results in agreement with the CMB dipole albeit with large error bars.
@@
In previous work [22] we determined the dipole inferred from the Pantheon+ compilation of type Ia Supernovae [23]. We found a dipole compatible in amplitude with the CMB dipole, but pointing in a different direction. However, the dipole amplitude in supernova distances is proportional to $[r(z)H(z)]^{-1}$ and hence it rapidly decays with redshift so that for supernovae with $z>0.1$ no significant dipole can be measured. It is therefore possible that the dipole we have seen in this data actually corresponds to the velocity ${\mathbf{v}}_{\odot}-{\mathbf{v}}^{\rm(bulk)}$ where ${\mathbf{v}}_{\odot}$ is the peculiar velocity of the solar system and ${\mathbf{v}}^{\rm(bulk)}$ is the bulk velocity of a sphere around us with radius $R\lesssim z/H_{0}=300(z/0.1)h^{-1}$pc, where $H_{0}$ is the present value of the Hubble parameter. Interestingly, the bulk velocity inferred in this way is in relatively good agreement with the result of the ‘cosmicflow4’ analysis [24], however our error-bars are significantly larger. In the present paper we test this hypothesis. As we discuss in the next section, if the SN dipole is really due to the peculiar motion of the individual supernovae and not due to a global dipole, we expect to observe also a monopole and a quadrupole (and higher multipoles) of similar amplitude. For this reason, we determine in this paper also the monopole and the quadrupole of the Pantheon+ compilation of supernova distances, in addition to the bulk flow (dipole) – a possible quadrupolar Hubble expansion in Pantheon+ was also studied in [25] with the help of a cosmographic expansion. We do indeed find a monopole and a quadrupole with amplitudes of the expected order of magnitude. We also argue that this amplitude for the bulk velocity is not extremely unlikely in the standard $\Lambda$CDM model.
@@ 2 Theoretical description
Here ${\mathbf{v}}_{\odot}$ is the observer velocity, $r(z)$ is the comoving distance out to redshift $z$, ${\mathbf{v}}({\mathbf{n}},z)$ is the peculiar velocity of the source. The functions are to be evaluated at ${\mathbf{x}}={\mathbf{n}}r(z)$ and $t=t(z)=t_{0}-r(z)$. The symbol $\Delta_{\Omega}$ denotes the Laplacian on the sphere, while $\bar{d}_{L}$ is the luminosity distance of the background FL universe. In a flat $\Lambda$CDM universe at low redshift where radiation can be neglected it is given by
@@ 2.1 The monopole and quadrupole perturbations of the luminosity distance at low redshift
If the source peculiar velocity ${\mathbf{v}}({\mathbf{n}},z)$ were independent of direction, a pure ‘bulk velocity’ shared by all supernovae, this would lead to a pure dipole, which is what we considered in our previous paper. However, we expect that there should also be some dependence on direction and on redshift. Here we assume that the redshift dependence is the one given by linear perturbation theory, this is a reasonable assumption for the large scales that we investigate (e.g. $z=0.025$ corresponds to a radius of $75h^{-1}$pc). We then fit for an angular dependence in the form of a monopole and a quadrupole. Of course we expect in principle also higher multipoles to be present but we neglect them here. As the different multipoles are orthogonal to each other, this should not bias our results on the monopole, dipole and quadrupole. Within linear perturbation theory, the time dependence of the peculiar velocity field is given by
@@
In this expression, $\alpha_{ij}$ is a symmetric traceless tensor which represents the ‘bulk quadrupole’ today, and since $n^{i}\delta_{ij}n^{j}=1$, $\gamma$ corresponds to a monopole, the part of ${\mathbf{v}}$ that is parallel to the radial direction ${\mathbf{n}}$. Both, $\alpha_{ij}$ and $\gamma$ have the units of a velocity. Putting all of this together we obtain
@@
In Fig. 1 we show the behaviour of the prefactor $A(z)$. At small redshifts, $z\lesssim 0.2$, we find that $A(z)\simeq 1$, implying that the correction due to this factor is negligible, in agreement with the assumption of a bulk motion of nearby galaxies.
Figure 1: The prefactor A(z) as function of redshift. We assume $H_{0}=73.4$ km/s/Mpc and $\Omega_{{\rm m}}=0.338$.
@@ 2.2 Redshift corrections
$z^{\text{(cmb)}}(z,{\mathbf{n}})=\frac{1+z}{1+z_{\odot}}-1\,,$ (2.14)
@@
$z_{\odot}=\sqrt{\frac{1+(-v_{\odot})/c}{1-(-v_{\odot})/c}}-1\,,$ (2.15)
@@
$z_{q}(z,{\mathbf{n}})=\frac{1+z_{\odot}}{1+z_{p}(z,{\mathbf{n}})}-1=\frac{1+z}{(1+z^{\text{(cmb)}})(1+z_{p}(z,{\mathbf{n}}))}-1\,,$ (2.16)
@@
We can then rewrite (2.8) more concisely as
@@ 3 Data and methodology
$\mu=5\log_{10}(d_{L}/10{\rm pc})=5\log_{10}(d_{L}/1{\rm Mpc})+25=5\log_{10}d_{L}+M\,.$ (3.1)
As reported in Tab. 8, 77 SNe are in galaxies that also host Cepheids, for which we know the absolute distance modulus $\mu_{\mathrm{ceph}}$. While Pantheon+ uses corrected redshifts including the motion of the solar system and estimated peculiar velocities of the sources, we use the actually measured redshifts for our analysis.
For handling the astronomical quantities and convenient unit conversions we use the package astropy [32, 33, 34]. To determine the parameters of our model we perform an MCMC analysis using the python package emcee [35]. Our code is parallelized using the Python package schwimmbad [36]. Our sampler consists of 32 walkers with the “stretch move” ensemble method described in [37].
@@
where $C$ is the covariance matrix provided by the Pantheon+ collaboration. The vector $\Delta\bm{\mu}$ is defined by
@@
Table 1: Uniform priors for the parameters sampled in the various MCMC analyses. We vary $\sin(\text{dec})$ and not dec in order to sample the celestial sphere uniformly in area. We then apply the $\arcsin$ function to the chain entries in order to recover the declination for the analysis. $\alpha_{ij}$ refers to a generic element of the quadrupole matrix introduced in Equation (2.8).
@@ 4 Results
Figure 2: Applying the redshift correction as described in Section 2.2 (orange contours), we obtain the same results as in our previous paper [22] (blue contours) in which we applied a correction at the level of the definition of luminosity distance (Equation (2.8)). The dashed lines show as reference values the Planck dipole values in (1.1) for ${\mathbf{v}}_{\odot}$ (the Planck value for the declination is outside the plot range), dM$=0$ and the cosmological parameters $\Omega_{{\rm m}}=0.338$, $H_{0}=73.4$ km/s/Mpc given by Pantheon+ [23].
As we see in Figure 2, we obtain the same contours as in our previous paper where the corrections were applied at the level of the definition of luminosity distance instead of the redshift, i.e., using only the first term in the Taylor series in $\bar{d}_{L}(\bar{z}+\delta z)$ and linearizing $\delta z$ in ${\mathbf{n}}{\cdot}{\mathbf{v}}_{\odot}$ (neglecting bulk velocity, the monopole and the quadrupole). The robustness of applying redshift corrections is also manifest by the fact that, when considering the dipole only, we find a negligible difference of $\Delta\chi^{2}\simeq 0.8$ with respect to the analysis developed in the previous paper where we used Eq. (2.11). In Table 2 we report the constraints inferred from the MCMC routine for the new orange contours, where the bulk velocity correction is taken into account as a redshift correction. Note that while the amplitude roughly agrees with the velocity of the solar system inferred from the CMB, the direction is very different, compared with Eq. (1.1). This result is in excellent agreement with our previous paper [22], where it was the main finding. It led us to the conclusion that the bulk velocity cannot be neglected.
$|{\mathbf{v}}_{\odot}|$[km/s] ra[°] dec[°] dM $H_{0}$[km/s/Mpc] $\Omega_{{\rm m}}$
$318\pm 40$ $140\pm 7.7$ $42^{+7}_{-6}$ $-0.003\pm 0.029$ $73.2\pm 1.0$ $0.340\pm 0.018$
Table 2: Constraints on parameters for the dipole inferred in the Pantheon+ data set according to Equation (2.16) neglecting $v_{p}$ corrections. Here and in all the following results tables, the errors show the 68% confidence errors obtained analysing the MCMC routine with getdist. They are purely statistical errors, as a consequence, they should be interpreted with care.
4.1 Bulk velocity
When fixing ${\mathbf{v}}_{\odot}$ to the Planck value and including just the ${\mathbf{v}}^{(\rm bulk)}$ correction in Equation (2.17), i.e. considering only the first term in brackets in Equation (2.18), the dipole, we obtain the contour plots shown in Figure 3 and reported in Table 3. These bulk velocities agree with our previous paper [22].
@@ 4.1 Bulk velocity
$|{\mathbf{v}}^{(\rm bulk)}|$[km/s] ra(bulk)[°] dec(bulk)[°] dM $H_{0}$[km/s/Mpc] $\Omega_{{\rm m}}$
$320\pm 40$ $203\pm 11$ $-52.5^{+4.8}_{-5.5}$ $-0.001\pm 0.029$ $73.2\pm 1.0$ $0.340\pm 0.018$
Table 3: Constraints on parameters for the bulk motions inferred in the Pantheon+ data set according to Equation (2.16) neglecting quadrupole corrections and fixing ${\mathbf{v}}_{0}$ to the Planck value (1.1).
@@ 4.2 Bulk + quadrupole analysis
We now include also the quadrupole in the luminosity distance, which actually comes from the dipole in the peculiar velocity and which we describe by the matrix $\alpha$, to perform a bulk + quadrupole analysis. As mentioned in Section 2.1, $\alpha$ is a symmetric trace-less tensor of dimension $3\times 3$. It is defined by five parameters that we introduce in our MCMC routine.
@@
In our parametrisation of the matrix $(\alpha_{ij})$ the trace vanishes, hence the sum of the eigenvalues is zero by construction. However, eigenvalues $\lambda_{1}$ and $\lambda_{3}$ differ from zero by 3 to 4 standard deviations. Since their distributions are close to Gaussian, see Fig. 4, we conclude that the detection of the quadrupole is significant. The direction of the eigenvectors is fixed up to a sign. To remove this ambiguity, we choose all eigenvectors to point into the northern hemisphere. They are normalized and dimensionless, since we assume the eigenvalues to have the dimension of velocity. In Table 5 we report their directions. As $\lambda_{2}$ is compatible with zero, it is not surprising that the direction of ${\mathbf{w}}_{2}$ is not well determined. However, also the direction of ${\mathbf{w}}_{1}$ has surprisingly large errors.
@@
and compare it with $|{\mathbf{v}}^{(\rm bulk)}|$ obtained in the bulk + quadrupole analysis. While $|{\mathbf{v}}^{(\rm bulk)}|=338\pm 40\,$km/s, we find $\lambda=190\pm 40\,$km/s. [^1] Errors associated to $\lambda$ are computed using the python package uncertainties [39]. Even though the quadrupole is somewhat smaller than the dipole, it is of a comparable order of magnitude.
@@
[…]
In Fig. 5 and Table 5 we also show the scalar products $\mu_{i}^{2}=\left[({\mathbf{w}}_{i}\cdot{\mathbf{v}}^{\text{(bulk)}})/|{\mathbf{v}}^{\text{(bulk)}}|\right]^{2}$. Note that the sign has no significance since both ${\mathbf{w}}_{i}$ and $-{\mathbf{w}}_{i}$ are eigenvectors of $\lambda_{i}$. Since the direction of ${\mathbf{w}}_{2}$ is not well determined the value of $\mu_{2}$ is also not. Interestingly, ${\mathbf{w}}_{3}$ is well aligned with ${\mathbf{v}}^{\rm(bulk)}$ and, as a consequence, ${\mathbf{w}}_{1}$ is nearly orthogonal to ${\mathbf{v}}^{\rm(bulk)}$.
Figure 6: In this figure we show the parameters not involving the quadrupole and compare their values in an analyis including the quadrupole (orange) and an analysis containing only the dipole from the bulk velocity.
In Fig. 6 we compare the inferred cosmological parameters and the bulk velocity in an analysis including the quadrupole from the bulk motion (orange) with the one including only the dipole. We find that all values are virtually identical in both analyses. It is not surprising that the dipole is not changed, as we expect different multipoles to be independent, but we note that also the cosmological parameter constraints stay the same.

### v2

@@ Abstract
In previous work we have shown that the dipole in the low redshift supernovae of the Pantheon+SH0ES data does not agree with the one inferred from the velocity of the solar system as obtained from CMB data. We interpreted this as the presence of significant bulk velocities, indicating that it could be interesting to look at other large-scale multipoles. In this paper we study the monopole, dipole and quadrupole in the Pantheon+SH0ES data. We find that in addition to the dipole also both the monopole and the quadrupole are detected with high significance. They are of similar amplitudes as the bulk flow. While the monopole is only significant at very low redshift, the quadrupole even increases with redshift.
@@ 1 Introduction
Due to our motion with respect to the surface of last scattering, the CMB also exhibits a dipole with an amplitude of about $10^{-3}$. This dipole has been discovered in the 1970s [5, 6] and is now measured with exquisite precision [7, 8, 1]. This anisotropy in the CMB also leads to the correlation of adjacent multipoles which have consistently been measured with a significance of about 5 standard deviations [9]. Attributing the entire CMB dipole to our motion, one infers a velocity of the solar system given by
@@
where $({\rm ra,dec})$ are the ‘right ascension’ (ra) and ‘declination’ (dec) denoting the directions with respect to the barycenter of the solar system (at J2000, i.e. January 1, 2000). A possible intrinsic dipole in the CMB of the same order as the higher multipoles is expected to change this result by about 1%. (This is simply a consequence of the fact that the clustering $C_{\ell}$’s are about four orders of magnitude smaller than $C_{1}$. Assuming that $C_{1}$ has a similar intrinsic clustering contribution this changes the dipole $\propto\sqrt{C_{1}}$ by about 1%.)
@@
Within the standard model of cosmology we expect to see this dipole due to our motion also in the large scale distribution of galaxies [10]. While first results of a radio survey agreed reasonably well with the CMB velocity [11], more recent analyses of catalogs of radio galaxies and quasars have found widely differing results from which significantly larger peculiar velocities have been inferred [12, 13, 14, 15, 16]. The latest results [17] show a $5\sigma$ discrepancy with the CMB dipole which is considered by the authors as a challenge of the cosmological principle. There are, however also critiques that the analysis of the data might be too simplified and that a more refined analysis could give results that are consistent with standard cosmology [18, 19, 20], see also [21] for an alternative method which gives results in agreement with the CMB dipole albeit with large error bars.
@@
In previous work [22] we determined the dipole inferred from the Pantheon+ compilation of type Ia supernovae [23]. The first attempt to measure the dipole in supernova data dates back to 2006, however with a much smaller dataset and large error bars [24]. In [22] we found a dipole compatible in amplitude with the CMB dipole, but pointing in a different direction. However, the dipole amplitude in supernova distances is proportional to $[r(z)H(z)]^{-1}$ and hence it rapidly decays with redshift so that for supernovae with $z>0.1$ no significant dipole can be measured. Here $r(z)$ is the comoving distance out to redshift $z$ and $H(z)$ is the Hubble parameter at redshift $z$. It is therefore possible that the dipole we have seen in this data actually corresponds to the velocity ${\mathbf{v}}_{\odot}-{\mathbf{v}}^{\rm(bulk)}$ where ${\mathbf{v}}_{\odot}$ is the peculiar velocity of the solar system and ${\mathbf{v}}^{\rm(bulk)}$ is the bulk velocity of a sphere around us with radius $R\lesssim z/H_{0}=300(z/0.1)h^{-1}$pc, where $H_{0}$ is the present value of the Hubble parameter. Here we use the Hubble law for small redshifts, $z\ll 1$ : the radius out to $z$ is given by $R(z)=z/H_{0}$ and $H_{0}=h/(3000$Mpc)))). The speed of light is set to $c=1$ in our formulae. Interestingly, the bulk velocity inferred in this way is in relatively good agreement with the result of the CosmicFlows4 analysis [25], however our error-bars are significantly larger. In the present paper we test this hypothesis. As we discuss in the next section, if the SN dipole is really due to the peculiar motion of the individual supernovae and not due to a global dipole, we expect to observe also a monopole and a quadrupole (and higher multipoles) of similar amplitude. For this reason, we determine in this paper also the monopole and the quadrupole of the Pantheon+ compilation of supernova distances, in addition to the bulk flow (dipole) – a possible quadrupolar Hubble expansion in Pantheon+ was also studied in [26] with the help of a cosmographic expansion.
We do indeed find a monopole and a quadrupole with amplitudes of the expected order of magnitude. We also argue that the amplitude we find for the bulk velocity is not extremely unlikely in the standard $\Lambda$CDM model.
@@ 2 Theoretical description
Here ${\mathbf{v}}_{\odot}$ is the observer velocity and ${\mathbf{v}}({\mathbf{n}},z)$ is the peculiar velocity of the source. The functions are to be evaluated at ${\mathbf{x}}={\mathbf{n}}r(z)$ and $t=t(z)=t_{0}-r(z)$. The symbol $\Delta_{\Omega}$ denotes the Laplacian on the sphere, while $\bar{d}_{L}$ is the luminosity distance of the background FL universe. In (2.1) and in the following, the redshift $z$ is the observed, measured redshift. In the Pantheon+ data release this redshift is denoted $z{\rm HEL}$, indicating heliocentric redshift.
In a flat $\Lambda$CDM universe at low redshift where radiation can be neglected it is given by
@@ 2.1 The monopole and quadrupole perturbations of the luminosity distance at low redshift
If the source peculiar velocity ${\mathbf{v}}({\mathbf{n}},z)$ were independent of direction, a pure ‘bulk velocity’ shared by all supernovae, this would lead to a pure dipole, which is what we considered in our previous paper. However, we expect that the peculiar velocity also depends on direction and on redshift. Here we assume that the redshift dependence is the one given by linear perturbation theory, this is a reasonable assumption for the large scales that we investigate (e.g. $z=0.025$ corresponds to a radius of $75h^{-1}$pc). We then fit for an angular dependence in the form of a monopole and a quadrupole. Of course we expect in principle also higher multipoles to be present but we neglect them here. As the different multipoles are orthogonal to each other, this should not bias our results on the monopole, dipole and quadrupole [^1] If the sky coverage would be homogeneous we would expect the different multipoles to be completely independent of each other. But since this is not the case, some bias might still be present.. Within linear perturbation theory, the time dependence of the peculiar velocity field is given by
@@
In this expression, $\alpha_{ij}$ is a symmetric traceless tensor which can be given, e.g. by the five components $\alpha_{11}$, $\alpha_{22}$, $\alpha_{12}$, $\alpha_{13}$ and $\alpha_{23}$. It represents the quadrupole of the peculiar velocity field today, and since $n^{i}\delta_{ij}n^{j}=1$, $\gamma$ corresponds to a monopole, the part of ${\mathbf{v}}$ that is parallel to the radial direction ${\mathbf{n}}$. Both, $\alpha_{ij}$ and $\gamma$ have the units of a velocity. Putting all of this together we obtain
@@
In figure 1 we show the behaviour of the prefactor $A(z)$. At small redshifts, $z\lesssim 0.2$, we find that $A(z)\simeq 1$, implying that the correction due to this factor is negligible, in agreement with the assumption of a bulk motion of nearby galaxies.
Figure 1: The prefactor $A(z)$ as function of redshift. We assume $H_{0}=73.6$ km/s/Mpc and $\Omega_{{\rm m}}=0.334$.
@@
where ${}_{2}F_{1}(a,b;c;d)$ represents the confluent hypergeometric function (see [31], Chapter 13) and
@@ 2.2 Redshift corrections
$1+z^{\text{(cmb)}}(z,{\mathbf{n}})=\frac{1+z}{1+z_{\odot}}\,,$ (2.14)
@@
$1+z_{\odot}=\sqrt{\frac{1+(-v_{\odot})/c}{1-(-v_{\odot})/c}}\,,$ (2.15)
@@
$1+z_{q}(z,{\mathbf{n}})=\frac{1+z^{\text{(cmb)}}}{1+z_{p}(z,{\mathbf{n}})}=\frac{1+z}{(1+z_{\odot})(1+z_{p}(z,{\mathbf{n}}))}\,,$ (2.16)
@@
For (2.16) we simply use the formula for the redshift combining two boosts, one to the CMB frame and one due to the peculiar velocity of the source. We can then rewrite (2.8) more concisely as
@@ 3 Data and methodology
$\mu=5\log_{10}(d_{L}/10{\rm pc})=5\log_{10}(d_{L}/1{\rm Mpc})+25\,.$ (3.1)
As reported in table 6, 77 SNe are in galaxies that also host Cepheids, for which we know the absolute distance modulus $\mu_{\mathrm{ceph}}$. While Pantheon+ uses corrected redshifts including the motion of the solar system and estimated peculiar velocities of the sources, we use the actually measured redshifts for our analysis.
For handling the astronomical quantities and convenient unit conversions we use the package astropy [35, 36, 37]. The theoretical model introduced in section 2 is implemented in the code scoutpip [^2] Available at https://github.com/fsorrenti/scoutpip. To determine the parameters of our model we perform an MCMC analysis using the python package emcee [38]. Our code is parallelized using the Python package schwimmbad [39]. Our sampler consists of 32 walkers with the “stretch move” ensemble method described in [40].
@@
where $C$ is the covariance matrix provided by the Pantheon+ collaboration [^3] We use the full covariance matrix, the sum of statistical and systematic errors.. The vector $\Delta\bm{\mu}$ is defined by
@@
Table 1: Uniform priors for the parameters sampled in the various MCMC analyses. We vary $\sin(\text{dec})$ and not dec in order to sample the celestial sphere uniformly in area. We then apply the $\arcsin$ function to the chain entries in order to recover the declination for the analysis. $\alpha_{ij}$ refers to a generic element of the quadrupole matrix introduced in eq. (2.8).
@@ 4.1 Simple dipole analysis
Doing this we obtain the same results as in our previous paper where the corrections were applied at the level of the definition of luminosity distance instead of the redshift, i.e., using only the first term in the Taylor series in $\bar{d}_{L}(\bar{z}+\delta z)$ and linearising $\delta z$ in ${\mathbf{n}}{\cdot}{\mathbf{v}}_{\odot}$ (neglecting bulk velocity, the monopole and the quadrupole). The robustness of applying redshift corrections is also manifest by the fact that, when considering the dipole only, we find a negligible difference of $\Delta\chi^{2}\simeq 0.8$ with respect to the analysis developed in the previous paper where we used eq. (2.11). In the first row of table 4 we report the constraints inferred from the MCMC routine for the new orange contours, where the bulk velocity correction is taken into account as a redshift correction. Note that while the amplitude roughly agrees with the velocity of the solar system inferred from the CMB, the direction is very different, compared with eq. (1.1). This result is in excellent agreement with our previous paper [22], where it was the main finding. It led us to the conclusion that the bulk velocity cannot be neglected.
4.2 Including a bulk velocity
When fixing ${\mathbf{v}}_{\odot}$ to the Planck value and including just the ${\mathbf{v}}^{(\rm bulk)}$ correction in equation (2.17), i.e. considering only the first term in brackets in equation (2.18), the dipole, we obtain the contour plots shown in figure 2 and reported in the second row of table 4. These bulk velocities agree with our previous paper [22]. As discussed there in details, the direction of the fitted bulk flow agrees well with the bulk flow direction assumed in the Pantheon+ analysis [23], but the amplitude is nearly twice as large.
@@ 4.3 Bulk + quadrupole analysis
We now include also the quadrupole in the luminosity distance, which comes from the angular dependence of the peculiar velocity field as discussed in section 2.1 and which we describe by the matrix $(\alpha_{ij})$, to perform a bulk + quadrupole analysis. As mentioned in section 2.1, $(\alpha_{ij})$ is a symmetric trace-less tensor of dimension $3\times 3$. It is defined by five parameters (e.g. $\alpha_{11},~{}\alpha_{22},~{}\alpha_{12},~{}\alpha_{13}$ and $\alpha_{23}$) that we introduce in our MCMC routine. In Fig. 3 we visualise the quadrupole contribution plotting the function $Q({\mathbf{n}})=\alpha_{ij}n^{i}n^{j}$.
Figure 3: We show the inferred quadrupole function, $Q({\mathbf{n}})=\alpha_{ij}n^{i}n^{j}$ in equatorial coordinates. The Supernova positions as well as the direction of the bulk velocity are also indicated. Note that the peak of the quadrupole is not far from the bulk velocity direction. No strong correlation with the supernova positions is evident.
Instead of the matrix elements $\alpha_{ij}$, however, we show in our plots the geometrically more interesting quantities given by the eigenvalues and the direction of the eigenvectors. Note that these also amount to five parameters given e.g. by $\lambda_{1}$ and $\lambda_{2}$, the traceless condition then determines $\lambda_{3}$, as well as three angles determining the direction of the first two orthogonal eigenvectors. The third eigenvector is then simply given by the orthogonality condition, ${\mathbf{w}}_{3}={\mathbf{w}}_{1}\times{\mathbf{w}}_{2}$. Note that with ${\mathbf{w}}_{i}$ also $-{\mathbf{w}}_{i}$ is an eigenvector, so we can fix their orientation at will.
@@
Figure 5: The scalar product $\mu_{i}^{2}=\left[({\mathbf{w}}_{i}\cdot{\mathbf{v}}^{\text{(bulk)}})/|{\mathbf{v}}^{\text{(bulk)}}|\right]^{2}$, analysis without redshift cut.
@@
In our parametrisation of the matrix $(\alpha_{ij})$ the trace vanishes, hence the sum of the eigenvalues is zero by construction. However, eigenvalues $\lambda_{1}$ and $\lambda_{3}$ differ from zero by 3 to 4 standard deviations. Since their distributions are close to Gaussian, see figure 4, we conclude that the detection of the quadrupole is significant. The direction of the eigenvectors is fixed up to a sign. To remove this ambiguity, we choose all eigenvectors to point into the northern hemisphere. They are normalized and dimensionless, since we assume the eigenvalues to have the dimension of velocity. In table 3 we report their directions. As $\lambda_{2}$ is compatible with zero, it is not surprising that the direction of ${\mathbf{w}}_{2}$ is not well determined. However, also the direction of ${\mathbf{w}}_{1}$ has surprisingly large errors.
@@
and compare it with $|{\mathbf{v}}^{(\rm bulk)}|$ obtained in the bulk + quadrupole analysis. While $|{\mathbf{v}}^{(\rm bulk)}|=338\pm 40\,$km/s, we find $\lambda=190\pm 40\,$km/s. [^4] Errors associated to $\lambda$ are computed using the python package uncertainties [42], which handles error propagation to determine the errors of functions of quantities with uncertainties. Even though the quadrupole is somewhat smaller than the dipole, it is of a comparable order of magnitude.
@@
In figure 5 and table 3 we also show the scalar products $\mu_{i}^{2}=\left[({\mathbf{w}}_{i}\cdot{\mathbf{v}}^{\text{(bulk)}})/|{\mathbf{v}}^{\text{(bulk)}}|\right]^{2}$. Note that the sign has no significance since both ${\mathbf{w}}_{i}$ and $-{\mathbf{w}}_{i}$ are eigenvectors of $\lambda_{i}$. Since the direction of ${\mathbf{w}}_{2}$ is not well determined the value of $\mu_{2}$ is also not. Interestingly, ${\mathbf{w}}_{3}$ is well aligned with ${\mathbf{v}}^{\rm(bulk)}$ and, as a consequence, ${\mathbf{w}}_{1}$ is nearly orthogonal to ${\mathbf{v}}^{\rm(bulk)}$.
Figure 6: In this figure we show the parameters not involving the quadrupole and the monopole and compare their values in an analysis including the quadrupole (orange) and an analysis containing only the dipole from the bulk velocity.
In figure 6 we compare the inferred cosmological parameters and the bulk velocity in an analysis including the quadrupole from the bulk motion (orange) with the one including only the dipole. We find that all values are virtually identical in both analyses. It is not surprising that the dipole is not changed, as we expect different multipoles to be independent, but we note that also the cosmological parameter constraints stay the same.